25.00 g of potassium hydrogen phosphate (K HPO4) is dissolved in enough water to make 125.0 mL of a solution with a density of 1.22 g/mL. What is the concentration of the solution in % (m/m). % (m/v). molarity, mosM and mEq/L of the potassium ion? Note: K2HPO4 has a molar mass of 174.2g/mole. What would be the final molarity of the above solution if you were to dilute it by adding 175 mL of water?
Answer #1

25 g of K2HPO4

Total volume of solution =125 ml

Density of solution =1.22 g/ml

So total mass of solution = density *volume

= 152.5 g

Out of 152.5 g , 25 g is of KHPO4, so mass of water = 152.5-25 = 127.5 g

So %m/m = (25/152.5)*100 = 16.39%

%w/v = (25/125)*100= 20%

Number of moles of KHPO4 = (25/174.2) = 0.1435 mol

Molarity =number of moles /volume in litre

M1 = (0.1435/125)*1000 = 1.15 mol/L

Molarity of k+ = 2*1.15 = 2.3 mol/L

Now the solution is diluted by adding 175 ml water so total volume at this stage = 125+175=300ml

So M1V1= M2V2

(1.15*125) = M2*300

M2= 0.479 mol/L